OpenStudy (anonymous):

Write a rule that represents the function (0,2), (1,5), (2,9), (3,13), (4,17)

OpenStudy (pawanyadav):

It is (0,1) or (0,2) check again.

OpenStudy (anonymous):

(0,1)

OpenStudy (pawanyadav):

Yes, then it follows. y=4x+1 as 1=0*4+1 5=1*4+1 9=2*4+1 13=3*4+1 17=4*4+1